### 27 Mar 2026 v2

# Raman phonon dynamics and its control for enhanced optical frequency conversion

Abstract
Raman phonons arise from the inelastic scattering of light and represent quantized molecular motions that mediate a wide range of spectroscopic and nonlinear optical phenomena. In this work, we clarify the physical role of Raman phonons within a previously-developed time-domain framework based on the Raman-induced index modulation , and show that phonons correspond to the oscillatory component of the Raman-induced index modulation. The analysis further reveals a linear phonon-mediated interaction embedded within Raman scattering, in which optical fields couple through wave-vector matching with existing phonons. This mechanism underlies what has long been described as coherent Stokes and anti-Stokes scattering, as well as molecular modulation. Building on this insight, we introduce a phonon-controlled approach that enables efficient conversion into a selected Stokes order by tuning the wave-vector-matching relation between the driven phonons and the targeted Raman process. These results provide a clearer physical interpretation of Raman phonons and its corresponding Raman dynamics and offer new strategies for controlling Raman interactions.
1 Introduction
Raman phonon represents the quantized motions in matter that originate from the inelastic Raman scattering of light [1, 2, 3, 4, 5, 6, 7]. The scattered light is redshifted through exchange of energy between photons and internal degrees of freedom of the medium, driving molecular motions, and thus inducing phonons. This Raman phonon excitation provides direct access to the microscopic dynamics of matter and forms the basis of a wide range of spectroscopic [8, 9] and nonlinear optical phenomena [10, 11, 12].
Recently, we have developed a new perspective for understanding Raman dynamics based on the temporally-varying Raman-induced index modulation (Fig. 1) [13, 14], where is the Raman temporal response and is the optical field assuming an unchanged spatial profile, e.g., in an optical fiber. The nonlinear Raman field increment follows . Traditional approaches isolate only the oscillatory component of the index response by solving the coupled phonon equation with an imposed wave solution [: phonon wave vector; : phonon oscillation angular frequency which is equal to the Raman transition angular frequency] [15, 16, 17]. In contrast, this framework enables direct visualization of the full spectrotemporal Raman dynamics that combines frequency-dependent Raman gain spectrum and time-dependent phonon dynamics. In the long-pulse regime, the pulse-following component of contributes to Raman-enhanced self-phase modulation, facilitating supercontinuum generation and ultrashort-pulse formation [18, 19, 20, 21]. In the ultrashort-pulse regime, the pulse impulsively drives the Raman scattering whose inertial induces a significantly-delayed index modulation such that the pulse interacts with only the initial rising edge of the index wave. Operation in the intermediate regime combines features of both limits, yielding enhanced spectral broadening accompanied by a net redshift [22, 23, 24]. Moreover, the slow inertial response smooths the nonlinear index profile, facilitating high-fidelity spectral compression [25]. Although this framework provides a clear and comprehensive understanding of Raman dynamics across different temporal regimes, it leads to a new question regarding the connection between the index modulation and the traditional concept of quantized phonon. Is pulse-following index that also originates from Raman scattering a part of the Raman phonon? This article will explains the role of Raman phonon in terms of index modulation.
In addition to providing a fresh time-domain viewpoint, this framework reveals the linear phonon-mediated process embedded within Raman scattering. In a two-pulse configuration, phonons excited by the first pulse scatter the second pulse. Depending on the wave-vector matching between the pre-existing phonons and the Raman process driven by the second pulse, the scattering can produce Stokes [26, 27, 28, 29, 30] or anti-Stokes [31] radiation, corresponding to phonon generation or annihilation, respectively. Although this mechanism has been demonstrated experimentally, only recently have we provided a theoretical investigation [32, 13], and even then the discussion remained limited to two-pulse scenarios. Here, we further recognize that this linear phonon-mediated process constitutes the underlying origin of coherent Stokes and anti-Stokes Raman scattering (CSRS/CARS) [33, 34, 35, 36, 37, 38, 39, 40], whose well-known phase-matching condition originates from the phonon-mediated wave-vector matching. Moreover, this linear mechanism is distinct from the traditional interpretation based on nonlinear Raman (exponential) amplification. Due to the limitations of the traditional steady-state treatment, it has been historically overlooked. Notably, elucidating the underlying phonon interaction enables control of Raman generation beyond current capabilities. We introduce a phonon-controlled technique that suppresses early-stage cascaded Raman generation. By deliberately manipulating the light-phonon interaction, phonons generated by the first-order Stokes process exert minimal influence on subsequent Stokes generation, effectively confining the energy only to a specific order.
2 Raman phonon
Optical phonons have been well studied in crystalline solids using periodic Bloch-wave descriptions [41], where the relative motion of atoms within the basis gives rise to high-frequency vibrational modes that couple strongly to electromagnetic fields. These modes play a central role in infrared absorption, Raman scattering, and a wide range of light-matter interactions, owing to their ability to modulate the crystal’s polarization at optical frequencies.
The concept of optical phonon can be generalized beyond crystalline solids in the context of Raman scattering. The characteristic out-of-phase motion of atoms in a lattice provides a natural analogy to the vibrational dynamics of isolated molecules, allowing the optical-phonon picture to extend to fluids, i.e., liquids and gases, where no long-range periodic structure exists. In fact, the extension to fluids requires the ensemble-averaging over microscopic motions, and the collective oscillations of phonons refer to the macroscopic motions experienced by the optical field. This explains why a rotational Stokes field remains linearly polarized when driven by a linearly-polarized pump [42, 43, 24]. A linearly-polarized pump excites molecular rotations microscopically equally in the clockwise and counterclockwise directions, producing a macroscopic index modulation [Fig. 1(b)] that generates a linearly-polarized Stokes field. In contrast, a non-linearly-polarized driving field preferentially excites rotational Raman scattering in one direction, leading to a vectorial response in the resulting Stokes generation [13].
The internal motions of fluids involve vibrations and rotations of molecules, as well as intermolecular interactions. In gases, intermolecular interactions are typically weak and can often be neglected. Conventional descriptions of vibrational and rotational Raman dynamics, therefore, rely on individual molecular vibrations modeled as harmonic oscillators [17, 44, 45, 46] and molecular rotations modeled in terms of quantized rotational states [13, 47, 48, 49, 50], respectively (Fig. 2). The characteristic motions in these two models are indicated by the intramolecular displacement or molecular alignment angle between the light polarization (vector) and molecular axis . A more general framework employs the quantum-mechanical density matrix [51] to describe the molecular ensemble of both motions, and Raman dynamics in perturbative Raman operations, where the population remains predominantly in the ground state, are governed primarily by the off-diagonal coherence waves () [13]:
Connections between these distinct models can be established as follows [16, 13]:
with the reduced mass of a diatomic molecule, the vibrational transition angular frequency, the unperturbed population at level , the polarizability anisotropy, the real-valued electric field (with complex-valued representing its analytic signal), and the dephasing rate and the rotational transition angular frequency between rotational levels and , respectively. The coherence corresponds to the negative-frequency component of the displacement in vibrations, or directly proportional to the orientation in rotations, which, for example under linear polarization, becomes .
Phonons provide a quantum-mechanical description of the oscillatory component of Raman-induced index modulation (Fig. 3), namely only those oscillating around . In the impulsive regime, the induced index modulation exhibits a pronounced oscillatory behavior that persists beyond the duration of the driving pulse, a response that intuitively corresponds to the excitation of phonons. By contrast, in the long-pulse transient and steady-state regimes, the index predominantly follows the temporal envelope of the optical pulse and does not exhibit intrinsic oscillations. As pulse-following index does not induce an overall redshift for the pump but enhances spectrally-symmetric self-phase modulation, no net energy is transferred to the medium and thus no phonon is generated. Oscillatory behavior emerges only after Raman fields are generated, whose interference produces high-frequency beating that impulsively drives the oscillatory index response. The amplitude of these oscillations is determined by the strength of the optical beating, which in turn depends on the Stokes-field amplitude and its photon number. As one Stokes photon generated corresponds to one phonon generated, the magnitude of the oscillatory response reflects the number of phonons generated in the Raman process.
3 Linear phonon-mediated process
Traditional frameworks of Raman scattering focus primarily on the evolution of the optical fields, while the coupled phonon equation is treated more as an abstract auxiliary mathematical effect from vibrational [16, 31, 15, 52, 53, 54] or rotational [47, 48, 49, 50]. A clear physically-intuitive picture is obscured by having a coupled equation. In this article, we analyze Raman phenomena through the lens of the oscillatory Raman-induced index modulation that the field directly interacts with [13, 14], which makes the role of the phonon explicit and physically transparent. Within this framework, we delineate the interaction between phonons and optical fields, and elucidate the overlooked linear phonon-mediated process underlying Raman scattering.
Distinct nonlinear Raman phenomena emerge when an optical field propagates in a medium containing a pre-existing population of phonons. In the impulsive regime with an ultrashort driving field, the interaction reduces to the familiar femtosecond pump-probe configuration: depending on the relative temporal delay, the pulse can be redshifted, blueshifted, or temporally compressed [55, 56, 57, 58, 59, 60]. In contrast, the dynamics in the long-pulse transient and steady-state regimes are governed by wave-vector matching between the existing phonons and the Raman interaction [31, 32, 13]. Because newly-generated (absorbed) phonons are inherently phase-matched to the Stokes (anti-Stokes) process that creates (annihilates) them, this condition can equivalently be interpreted as wave-vector matching between the existing phonons and the phonons to be generated (or absorbed). Such wave-vector matching plays a crucial role in multi-field Raman interactions, for example in two-pulse schemes employing different wavelengths [32], or in cascaded Stokes processes of different orders that generate phonons having different wave vectors. This process follows (Fig. 4; see Supplementary Sec. 10 in [13] for derivation details):
where is the angular frequency of the [Stokes (S) or anti-Stokes (AS)] wave, , is the effective refractive index of the propagating mode, is the effective mode area, is the difference of the propagation constants of the wave and the pump, and the phonon wave follows . Unlike the nonlinear Raman amplification governed by the Raman gain, this phonon-mediated Raman generation is a linear process: it is linearly proportional to the pump field and the amplitude of the phonon waves . Because of its linear nature, its generation does not rely on any existing signals to amplify, but directly transfers energy from pump to Raman fields. Therefore, it is an efficient and thresholdless Raman generation mechanism during spontaneous Raman processes [61], which is particularly useful in seeding the latter nonlinear Raman amplification. In addition, the linear proportionality to shows that the generated Raman fields not only are coherent but also inherit the phase structure of the pump pulse. If the seed phonons have a wave vector that matches that of the generated phonons, [Eq. (3a)], seed phonons can be efficiently amplified, inducing optical Stokes generation. On the other hand, if the seed phonons have a wave vector that can be absorbed while satisfying the conservation of wave vector, and thus [Eq. (3b)], absorption occurs most efficiently and induces anti-Stokes generation.
This linear phonon-mediated process provides a complementary description of light-phonon interactions in terms of wave vector/momentum, supplementing the conventional energy-based picture given by and . Phonon absorption appears straightforward, as the associated wave-vector conservation naturally corresponds to the energy relation . However, a physically-intuitive description of phonon amplification is most naturally formulated in terms of phonon dynamics themselves, wherein amplification occurs most efficiently when the generated phonons closely resemble the seed phonons.
This process has long been considered as a four-wave-mixing (FWM) process, as in CSRS [33, 34, 35, 36] and CARS [37, 38, 39, 40]. In the traditional steady-state treatment, Raman temporal dynamics can be ignored and Raman-induced index modulation follows the pulse’s temporal profile, resembling the instantaneous electronic nonlinearity. In essence, the linear phonon-mediated process does involve four fields, two fields that produce the phonons and two fields in the latter phonon amplification or absorption process. Recent investigations in have demonstrated the importance of non-negligible temporal phonon dynamics due to the long-lived phonons [62, 63]. In particular, the probe pulse in CARS cannot be too strong; otherwise, its self-induced nonlinear Stokes generation can dominate over the linear anti-Stokes generation [as shown in Fig. 10(d) in [13]]. This behavior is difficult to reconcile with a description based solely on the nonlinear FWM that also exponentially amplifies the field [64]. As a result, instantaneous FWM is an inaccurate description of this linear phonon-mediated process, and therefore also of the CSRS and CARS.
Raman gain suppression occurs when the wave-vector-matching condition is satisfied due to a delicate balance between phonon annihilation and creation, yielding zero nonlinear Raman gain [15, 65, 66, 67, 17]. It originates from the coupled action between instantaneous nonlinear phase modulation and Raman gain, and therefore cannot be neglected even in media with strong Raman nonlinearity, where the instantaneous contribution is dominated by the pulse-following Raman-induced index in the long-pulse regime [13]. In contrast, for linear phonon-mediated interactions, the same condition corresponds to simultaneous generation of both Stokes and anti-Stokes fields [Eq. (3)]. Pre-existing phonons, generated by a strong pump and its Stokes field, seed and sustain the scattering processes, overcoming the nonlinear Raman gain suppression. This is precisely the mechanism that enables degenerate CARS under perfect wave-vector matching.
Similar to CSRS and CARS, molecular (phase) modulation is not a distinct mechanism but another manifestation of the linear phonon-mediated process. It represents a technique that modulates the incident field with pre-exciting coherent molecular motions, namely phonons. This approach has traditionally been employed for generating few-cycle pulses and frequency combs [68, 69, 70, 71, 72], or for frequency upconversion [31, 73, 74, 62]. The former arises from thresholdless coherent Raman generation of multiple orders, and the latter follows directly from Eq. (3b). Eq. (3) is the physically-intuitive counterpart to the recent quantum description of molecular modulation in [75]; however, the quantum treatment relies on the steady-state assumption that neglects phonon temporal dynamics. In contrast, our formulation [Eq. (3)] remains valid across timescales, including nonlocal interactions between two temporally-delayed pulses, which we will discuss in the next section. Quantum state preservation and entanglement arise from the linear proportionality to both the phonon wave and the converting pump field [Eq. (3)], which enforces a coherent correlation among all participating fields [62, 63, 75].
This process also underlies the coherent Stokes generation from spontaneous Raman scattering in the transient regime, where the phonon lifetime is much longer than the pulse duration (Fig. 5) [76]. To investigate it, we simulated with the unidirectional pulse propagation equation (UPPE) that considers realistic Raman response function for several types of gases [13], and the result is shown in Fig. 5. Initially, the Stokes field is seeded by vacuum fluctuations [77]. Due to the incoherence of vacuum fluctuations, the beating with the pump field remains fluctuating, leading to incoherent phonon response. The Stokes increment during propagation [14] is therefore dominated by the pump field (top figure in Fig. 5). However, limited phonon response time (due to the Raman period) results in a smoothed phonon wave with increased coherence that further drives more coherent Stokes generation through the linear phonon-mediated process. This new coherent Stokes wave overrides the initial incoherent Stokes wave, seeding the latter nonlinear amplification and eventually producing a coherent Stokes field that inherits the temporal phase from the pump [Eq. (3a)]. Moreover, the growing phonon wave leads to a Stokes increment that is delayed to the trailing edge of the pump pulse. Traditional frameworks based on nonlinear Raman amplification, in which the incoherent Stokes background is amplified, cannot adequately account for the emergence of coherence in this process. et al.Carman analyzed transient Stokes generation by assuming a priori that the Stokes and pump fields share the same phase structure as they evolve in a high-gain environment [76]. Only with the introduction of the linear phonon-mediated dynamics is this coherent generation now clearly explained.
Eq. (3) is introduced as an abstraction layer that isolates the linear phonon-mediated contribution from the full Raman response , clarifying the specific structural role this pathway plays in Raman temporal dynamics. Because the phonon wave is continuously driven by the optical fields, its amplitude and phase are not fixed material parameters but evolve with propagation distance. This construction therefore does not constitute an independent evolution equation; the actual dynamics must still be obtained from the full Raman formulation where explicitly contains both the field components that excite the phonons and those that are to be converted. The newly-introduced equations thus complement, rather than supplant, the conventional nonlinear description by selectively filtering out the linear phonon-mediated pathway to make its physical role more transparent.
4 Phonon-controlled Raman scattering
Linear phonon-mediated process underlies not only coherent transient Raman scattering but also early cascaded Raman generation. Fields of different Stokes orders can act as a new pump and interact with existing phonons generated by the first-order Stokes scattering, which simultaneously drives thresholdless Raman generation of various orders through linear phonon-mediated processes. Simultaneous growth of multiple Raman sidebands leads to reduced energy into a specific order. Most critically, continuous bidirectional nonlinear coupling among multiple Raman signals degrades the temporal fidelity of each individual signal. Current suppression techniques involve controlling Raman gain suppression by phase-matching Stokes and anti-Stokes fields [15], adding loss to the Raman band [78], inhibiting Raman feedback in a cavity [79], or constructing competition with modulation instability [80]; however, they completely suppress the Raman process. How to cleanly suppress Raman generation of only higher orders, or even target a specific order, remains unknown.
To mitigate this phonon-induced degradation, it is essential to sufficiently suppress the phonon-mediated Raman generation. When the wave-vector mismatch is nonzero [Eq. (3)], the process undergoes enhanced destructive interference over extended propagation distances, provided that the Raman nonlinearity is reduced accordingly (for example, by operating in a lower-pressure gas). Stronger destructive interference can also be achieved by directly increasing the wave-vector mismatch, which may be implemented by adjusting the gas pressure, introducing a buffer gas, or tailoring the hollow-core fiber structure.
Phonon-mediated processes depend on the wave-vector matching between the phonon wave vector and the corresponding Raman generation. To investigate its effect on early-stage cascaded Raman scattering, we limit the discussion to the second-order Stokes generation as an initialization of a cascaded process. To drive the second-order Stokes process with phonons generated from the first-order process, the phonon wave vector should match that from the second-order, that is, [Eq. (3a)], where represents the th-order Stokes field. Because , this cascaded wave-vector-matching relation is equivalent to the conventional wave-vector-matching relation but treating the first-order Stokes wavelength as a new pump:
In the following discussions, we will circulate around this for suppression of phonon-mediated processes.
To verify the idea of suppressing phonon-mediated processes, we simulated the vibrational Raman generation in -filled hollow-core fiber with the aforementioned UPPE. Fiber and pulse parameters followed [42] that generates a frequency comb. A linearly-polarized input pulse was seeded by a weak Stokes pulse. Despite seeding, it undergoes significant early-stage thresholdless cascaded vibrational Raman scattering due to small wave-vector mismatch with first-order phonons across multiple Stokes orders (partially indicated by a small here) [Fig. 6(a)]. Stokes pulses not only experience reduced energy but also suffer degraded temporal quality, manifested as pronounced temporal substructures resulting from energy exchange among the various Stokes orders. In addition, we adjusted the gas pressure to yielding and thereby enabling efficient energy transfer from the first to the second Stokes order [Fig. 6(b)]. Unexpectedly, the increased wave-vector mismatch for the phonon-mediated second- to third-order process () suppresses further cascaded growth, causing Raman generation to terminate at the second Stokes order and resulting in highly-efficient second-order Stokes conversion.
Figs. 6(c) and (d) illustrate complete suppression of phonon-mediated processes through an increased wave-vector mismatch over multiple orders [with only the first- to second-order displayed; Eq. (4)]. The enhanced wave-vector mismatch induces strong destructive interference, thereby inhibiting the phonon-mediated Raman generation and preventing the simultaneous growth of multiple Stokes orders. Under these conditions, the Raman process becomes effectively sequential: the Stokes signal of a given order grows only after the preceding order has reached its maximum, and this pattern repeats for each successive order. A Raman pulse that simultaneously achieves high energy and high temporal fidelity can thus be generated.
To assess its validity in the steady-state regime, we artificially reduced the Raman dephasing times while keeping the propagation distances fixed, and the resulting dynamics remain essentially unchanged (Fig. 7). This demonstrates that even in the steady-state limit where phonon temporal dynamics might appear negligible, the evolution is intrinsically transient: each temporal slice of the field is always influenced by phonons generated by earlier slices. More broadly, the universal action of the linear phonon-mediated process across all temporal regimes underscores that, irrespective of operating conditions, Raman scattering is fundamentally an ultrafast phonon-driven phenomenon governed by femtosecond phonon dynamics.
Moreover, we have verified this behavior using the two-pulse approach, which is widely employed to suppress Raman spectral narrowing (see Supplementary Fig. S2) [26, 27, 28, 29, 30]. When a chirped input pulse undergoes transient Raman generation, the Raman gain is strongest near its trailing edge [13], so only the trailing portion of the pulse is efficiently converted. This yields a Stokes pulse with a bandwidth narrower than that of the pump. In the two-pulse configuration, however, the phonon wave generated by the first pulse can linearly drive Raman generation across the entire duration of the chirped second pulse [Eq. (3)], thereby preventing Raman spectral narrowing. In the conditions demonstrated here, the phonons generated by the first pulse exhibit only small wave-vector mismatch with the second-order Stokes process driven by the second pulse. As a result, both the first- and second-order Stokes components grow simultaneously, giving rise to undesired early cascaded Raman generation. An increased wave-vector mismatch suppresses the cascaded process similarly. Most importantly, it shows that the two-pulse approach does not consistently yield a good first-order Stokes pulse. The effectiveness of this approach is tightly constrained by the wave-vector-matching conditions between phonons and subsequent Raman processes, which ultimately determine whether a phonon-mediated process is enhanced or suppressed.
5 Conclusion
In conclusion, we have illuminated the role of the Raman phonon through the unifying framework of the Raman-induced index modulation , spanning different temporal regimes, different media, and distinct Raman interaction types (see Supplementary Fig. S1 for a concluding figure). We further elucidated the linear phonon-mediated process, providing a complementary wave-vector-based perspective alongside the traditional energy-based description. This framework clarifies the underlying mechanisms of CSRS ,CARS, molecular modulation, and coherent spontaneous transient Raman generation. Finally, we introduced a phonon-controlled approach that deliberately enhances or inhibits the linear phonon-mediated process, enabling the suppression of early-stage cascaded Raman scattering and numerically achieving highly-efficient Stokes generation only at a specific order. The non-trivial contribution of phonons across distinct timescales underscores the inherently ultrafast time-dependent nature of Raman scattering, an aspect fundamentally missed by the traditional instantaneous steady-state assumption.
Funding National Institutes of Health (01EB033179, U01NS128660), the Office of Naval Research (N00014-19-1-2592), the Army Research Office (W911NF2410008), the Air Force Research Laboratory (FA8650-19-F-1025).
Acknowledgments Y.-H.C. was partially supported by a Mong Fellowship from Cornell Neurotech.
Disclosures The authors declare no conflicts of interest.
Data availability The code and the data used in this article have been made publicly available at https://github.com/AaHaHaa/gas_UPPE.